OpenStudy (anonymous):

does the following infinite series converge or diverge? explain your answer 7 - 21 +63 - 189 +...

OpenStudy (raden):

ratio = a2/a1 = -21/7 = -3 what kind is it ?

OpenStudy (amistre64):

it a big one :)